OpenStudy (anonymous):

Solve: 3x − x + 2 = 12

OpenStudy (ivettef365):

first combine like terms

OpenStudy (anonymous):

3x-x

OpenStudy (ivettef365):

exactly, so what is that

OpenStudy (anonymous):

3x

OpenStudy (ivettef365):

3x - x is the same as 3x - 1x so what is 3-1=?

OpenStudy (anonymous):

2

OpenStudy (uri):

3x − x + 2 = 12 2x+2=12 2x=12-2 Can you do it now? @doulo

OpenStudy (ivettef365):

ok , 2 so 3x - x = 2x so now you have 2x + 2 = 12

OpenStudy (ivettef365):

and uri is showing how to do the rest

hero (hero):

3x − x + 2 = 12 (3 - 1)x + 2 = 12 2x = 12 - 2 2x = 10 x = 10/2

OpenStudy (anonymous):

5 thanks

OpenStudy (ivettef365):

you got it